# SIGN: Safety-Aware Image-Goal Navigation for Autonomous Drones via Reinforcement Learning

### v1

@@ Abstract
Image-goal navigation (ImageNav) tasks a robot with autonomously exploring an unknown environment and reaching a location that visually matches a given target image. While prior works primarily study ImageNav for ground robots, enabling this capability for autonomous drones is substantially more challenging due to their need for high-frequency feedback control and global localization for stable flight. In this paper, we propose a novel sim-to-real framework that leverages visual reinforcement learning (RL) to achieve ImageNav for drones. To enhance visual representation ability, our approach trains the vision backbone with auxiliary tasks, including image perturbations and future transition prediction, which results in more effective policy training. The proposed algorithm enables end-to-end ImageNav with direct velocity control, eliminating the need for external localization. Furthermore, we integrate a depth-based safety module for real-time obstacle avoidance, allowing the drone to safely navigate in cluttered environments. Unlike most existing drone navigation methods that focus solely on reference tracking or obstacle avoidance, our framework supports comprehensive navigation behaviors–autonomous exploration, obstacle avoidance, and image-goal seeking–without requiring explicit global mapping. Code and model checkpoints will be released upon acceptance.
@@ I INTRODUCTION
Learning vision-based navigation has attracted increasing attention in autonomous drone research. Existing works [1, 2, 3] mostly focus on achieving autonomous obstacle avoidance and reference tracking, where the reference signal can be a target velocity [4], a direction [5], or goal coordinates [6]. Image navigation (ImageNav) poses a further challenge: enabling autonomous drones to search an unknown environment for a given goal image while avoiding collisions, a problem that remains largely unexplored. In this paper, we investigate ImageNav for autonomous drones using only visual and IMU measurements.
The body of ImageNav research has primarily focused on ground robots [7, 8, 9, 10, 11, 12]. Most existing studies are conducted exclusively in simulators such as Habitat [13] and neglect to penalize collisions with obstacles, which can result in unsafe behaviors when deployed in the real world [14]. Meanwhile, discrete action spaces such as {move_forward, turn_left, turn_right, stop} are commonly adopted to simplify learning by reducing the search space. This is generally only acceptable for ground robots, which are inherently stable and can correct their position through vision by the discrete low-frequency actions.
In contrast, achieving safe and efficient ImageNav for autonomous drones is considerably more challenging. Drones are inherently unstable platforms with fast dynamics, necessitating high-frequency feedback control and accurate global localization to maintain stable flight. When relying on visual–inertial localization, they are particularly susceptible to pose-estimation drift, especially in indoor environments. Furthermore, discrete action spaces often lack the precision and responsiveness needed for rapid pose correction. A continuous policy that directly outputs velocity commands can address this by enabling smoother and more timely control; however, it also expands the exploration space, substantially increasing training complexity. In addition, naively embedding safety constraints into policy learning can hinder effective exploration–particularly in cluttered or narrow environments such as those in Gibson [15], MP3D [16], and HM3D [17]–ultimately degrading navigation performance. Balancing efficient exploration with robust real-world safety therefore remains a central challenge in autonomous drone navigation.
In this paper, we introduce SIGN (Safety-aware Image-Goal Navigation), an end-to-end RL-based ImageNav framework for autonomous drones that employs continuous velocity control and incorporates enhanced safety mechanisms. Our main contributions are as follows:
@@ II-A Visual Navigation on Drones
We focus on end-to-end, map-less vision-based navigation without relying on global localization. Prior research in this direction has mainly addressed reference-tracking control with obstacle avoidance [1, 2, 3, 4, 5] or drone racing on a fixed course [18], where an RL policy maps sensor inputs directly to control commands such as velocity or acceleration. Similar to ImageNav tasks, [19] explored RL-based indoor search using drones, but their approach depends on global mapping. In contrast, our task is more challenging as it requires not only map-less exploration and collision avoidance, but also navigation toward a specified goal image.
@@ II-B End-to-end RL-based Visual Navigation
End-to-end RL has been shown to be highly efficient for ImageNav [9, 10, 11, 12], as it bypasses intermediate computations such as mapping and subgoal selection [8]. Early work by Wijmans [20] addressed point-goal visual indoor navigation for ground robots using a discrete action space; however, it relied on GPS and compass data, which is impractical for indoor settings. Al-Halah [9] proposed a map-less, end-to-end modular transfer learning framework for semantic visual navigation, enabling zero-shot generalization to new tasks and multi-modal goals (images, labels, or language) by aligning goal and observation embeddings in a shared latent space with a general-purpose semantic search policy. It is still limited to discrete actions, and pretraining–especially for multi-modal goal embeddings–can be computationally expensive depending on environmental diversity. Meanwhile, topological map-based navigation [21, 22, 23] remains effective by storing keyframes and leveraging graph neural networks to extract visual correlations for guidance, without relying on localization. Memory mechanisms [24] have also been incorporated to further enhance navigation performance.
@@ II-C Image-conditioned Waypoint Planner
Many works predict waypoints to reach a target image and use low-level position controllers for trajectory tracking [25, 26, 27], but they often rely on expert data and global localization to guide exploration. Imitation learning also suffers from out-of-distribution issues, especially when the current and target images have no overlap. In contrast, RL offers stronger exploration capabilities.
Figure 2: Performance comparison between discrete and continuous action using the SoTA baseline FGPrompt [10].
@@ III Visual Reinforcement Learning for ImageNav
In this section, we propose SIGN, a novel framework for drone ImageNav that leverages a continuous policy to enable map-less navigation with enhanced safety. Training continuous policies for ImageNav presents a significant challenge due to the expanded action space and increased training complexity. In Fig. 2, we compare a discrete-action State-of-The-Art (SoTA) baseline with its continuous counterpart, where the continuous action space $\mathcal{A}_{c}$ consists of forward linear and angular velocities $v_{lin},v_{ang}$. It can be observed that the discrete-action policy achieves steadily increasing rewards and much higher training efficiency, while the continuous-action policy suffers from very slow reward growth. This highlights the difficulty of training continuous policies in ImageNav tasks and motivates the design of our SIGN framework to address these challenges.
@@ III-B Navigation Model
Navigation model consists of a ResNet encoder $f_{\omega}$, a GRU model $g_{\xi}$, and an actor-critic policy trained by Proximal Policy Optimization (PPO) [28]. See Fig. 3 for the diagram. The goal image $O_{g}$ and current image $O_{c}$ are stacked as the input $s_{t}$. The inference procedure can be summarized as
@@ III-C Augmented Continuous Policy Training
A possible reason of the inefficient continuous policy training in Fig. 2 is that, in the early stage, training an effective vision encoder from scratch is time-consuming. RL policy struggles to improve until the vision encoder learns to capture key information from pixels. To address this problem, we propose two useful Self-Supervised Learning (SSL) techniques as auxiliary training tasks to enhance the policy learning.
Recall the standard optimization form in vanilla PPO [28]
@@ III-C1 Future Prediction
The prediction ability receives widespread attention due to its success in the model-based visual RL research [29, 30], where a foundation world model acts as an internal simulator to imagine future observations and guide policy learning. Focusing on a portable model for drone applications, here we train a single-step prediction to enhance the visual backbone training. To this end, we deploy an MLP head $P_{\psi}$ after the visual encoder $f_{\omega}$ to predict the reward $r_{t}$, the next-step latent feature $h_{t+1}=f_{\omega}(s_{t+1})$, and the done flag $d_{t}$, as illustrated in Fig. 4. The auxiliary training loss is defined as
@@
The forward prediction of dynamics $\hat{h}_{t+1}=P_{\psi}(a_{t},h_{t})$ is implemented in the latent space to avoid deconvolution operations and save computation resources. The reward loss function uses cross-entropy between the predicted reward $\hat{r}_{t}$ and a two-hot encoding of reward $r_{t}$, which proves more effective in predicting sparse rewards and robust to reward magnitude [30].
@@ III-C2 RandomShift
The RandomShift [31] technique performs data augmentation on the current image $O_{c}$ in each batch of samples. Denote the perturbated data as $aug(s_{t})=randomshift(O_{c})\oplus O_{g}$, then the auxiliary loss is defined as
@@
Input: number of env $B$, trajectory length $L$, fixed weights $\lambda_{r},\lambda_{d},\lambda_{T},\lambda_{rs}$
Initialize actor-critic models $A_{\theta}$ and $V_{\phi}$ including the backbone $f_{\omega}$ and GRU $g_{\xi}$
Initialize predictor model $P_{\psi}$ and collision predictor $Q_{c}$
for each epoch do
  Collect one trajectory $\{s_{t},a_{t},s_{t+1},r_{t},c_{t},d_{t}\}_{1:L}$ from each env, a total of $B\times L$ samples # Future Prediction
  for each mini-batch do
   Update $P_{\psi}$ and $f_{\omega}$ with data $\{s_{t},a_{t},s_{t+1},r_{t},d_{t}\}$ via minimizing Eq. (3)
  end for# PPO training with RandomShift
  for each mini-batch do
   Update policy parameters $\omega,\xi,\theta,\phi$ via Eq. (5)
   Train collision predictor $Q_{c}$ via minimizing Eq. (9)
  end for
end for
@@ IV Safety Module for Real-world Experiments
Existing ImageNav algorithms primarily focus on navigation performance while neglecting safety. Navigation policies trained in Habitat in those works are allowed to collide with obstacles. In our case, we also do not enforce safety during training, as doing so would greatly hinder exploration in narrow indoor environments. Instead, inspired by the safety layer designs in the SafeRL literature [32, 33], we propose a practical action correction mechanism that is activated only when the agent is in danger. In general, decoupling safety from navigation offers an effective strategy to simplify training and balance exploration with safety.
@@ IV-A Model Training for Collision Probability Prediction
Figure 5: Pipeline of the collision probability prediction.
The depth image is first preprocessed by cropping and min-pooling operations to obtain a depth vector $s^{d}$ ($s^{d}_{i}\in[0,1]$), indicating the closest distance in different directions. $Q_{c}$ is used to predict the collision probability based on the current depth image $s^{d}$ and navigation action $a_{t}$, defined as
@@
where $c_{t}$ is a collision feedback from the Habitat simulator, and the whole pipeline is shown in Fig. 5.
@@
$c_{t}^{\text{soft}}=\text{Sigmoid}\left(10\cdot(\beta-\min(s^{d}_{t}))\right),$ (7)
where $\beta$ is the threshold for a safe distance. When the minimum perceived distance falls below $\beta$, the soft label value increases rapidly toward 1, and it decreases to 0 when the distance exceeds $\beta$. Thus, collision detection is defined with respect to a safety margin distance of $\beta$.
@@ IV-B Action Correction for Obstacle Avoidance
The predicted collision probability is conditioned on $s_{t}^{d}$ and $a_{t}$, which means we can modify the action to reduce the collision probability. When the value of $Q_{c}$ exceeds a pre-defined threshold $d_{c}$, it indicates a dangerous situation, then actions $a_{t}$ from the navigation policy will be corrected iteratively until $Q_{c}<d_{c}$. The correction function $\psi(\cdot)$ will project the raw action to a safer one based on gradients [32]
$a_{k+1}=\psi(a_{k})=a_{k}-\frac{\partial}{\partial a_{k}}[Q_{c}(s^{d},a_{k})-d_{c}]^{+},$ (10)
where $k$ is the iteration number. The adjustment of the probability threshold $d_{c}$ can be viewed as a trade-off between safety and navigation performance. In general cases, $d_{c}$ is set to 0.5 because when $Q_{c}>0.5$, the minimum distance is less than the safety distance $\beta$ according to Eq. (7).
@@ V-A1 Datasets
Our navigation model is trained on Gibson ImageNav dataset [15], including 72 training scenes and 14 testing scenes. To further validate the robustness of our algorithm, the trained model is also evaluated on cross-domain datasets MP3D [16] and HM3D [17]. Scenes from different datasets have varying backgrounds and layouts but are all simulated in the Habitat.
@@ V-A2 Agent Configurations
To be consistent with SoTA works, we employ the Habitat camera that has a $90^{\circ}$ Field of View (FoV) and $128\times 128$ resolution. The continuous action space $\mathcal{A}_{c}$ consists of forward linear velocity $v_{lin}\in[0,0.25]~m/s$ and angular velocity $v_{ang}\in[-15^{\circ}/s,15^{\circ}/s]$. To implement the ‘Stop’ action in a continuous setting, the episode will be terminated early when $v_{lin}<0.025~m/s\ \&\ |v_{ang}|<1.5^{\circ}/s$.
V-A3 Reward and Performance Metrics
The design of the reward function $r_{t}$ follows ZER [9], which encourages the policy to approach the target and reduce the angle between the current view and the view of the goal.
@@ V-A3 Reward and Performance Metrics
Mem-Aug [24] ResNet18 4 RGB Discrete 56.0% 69.0%
VGM [21] ResNet18 4 RGB Discrete 64.0% 76.0%
TSGM [22] ResNet18 4 RGB Discrete 67.2% 81.1%
ZER [9] ResNet9 RGB Discrete 21.6% 29.2%
ZSON [12] ResNet50 RGB Discrete 28.0% 36.9%
OVRL-V2 [11] ViT-Base RGB Discrete 58.7% 82.0%
FGPrompt [10] ResNet9 RGB Discrete 66.5% 90.4%
FGPrompt [10] ResNet9 RGB Continuous 35.0% 79.9%
@@
Mem-Aug [24] Discrete $3.9\%$ $6.9\%$ $3.5\%$ $1.9\%$
NRNS [23] Discrete $5.2\%$ $9.3\%$ $4.3\%$ $6.6\%$
ZER [9] Discrete $10.8\%$ $14.6\%$ $6.3\%$ $9.6\%$
FGPrompt [10] Discrete $50.4\%$ $77.6\%$ $49.6\%$ $76.1\%$
FGPrompt [10] Continuous $22.9\%$ $59.4\%$ $20.1\%$ $56.5\%$
@@
In our experiments, $d_{s}=1m$ and $\alpha_{s}=25^{\circ}$ denote the success thresholds for distance and angular deviation, respectively. $r_{suc}=2.5$ is the success reward and $\gamma=-0.01$ is the slack penalty. On benchmarks, Success Rate (SR) and Success Weighted by Path Length (SPL) are two main metrics to quantify the navigation performance [34]. SPL is calculated as $\frac{1}{N}\Sigma^{N}_{i=1}S_{i}\frac{L_{i}}{max(L_{i},p_{i})}$, where $N$ is the total number of episodes, $S_{i}\in\{0,1\}$ is a success indicator, $p_{i}$ is the actual path length in the $i$-th episode, and $L_{i}$ is the optimal path length from the start to the goal point.

### v2

@@ Abstract
Image-goal navigation (ImageNav) tasks a robot with autonomously exploring an unknown environment and reaching a location that visually matches a given target image. While prior works primarily study ImageNav for ground robots, enabling this capability for autonomous drones is substantially more challenging due to their need for high-frequency feedback control and global localization for stable flight. In this paper, we propose a novel sim-to-real framework that leverages reinforcement learning (RL) to achieve ImageNav for drones. To enhance visual representation ability, our approach trains the vision backbone with auxiliary tasks, including image perturbations and future transition prediction, which results in more effective policy training. The proposed algorithm enables end-to-end ImageNav with direct velocity control, eliminating the need for external localization. Furthermore, we integrate a depth-based safety module for real-time obstacle avoidance, allowing the drone to safely navigate in cluttered environments. Unlike most existing drone navigation methods that focus solely on reference tracking or obstacle avoidance, our framework supports comprehensive navigation behaviors, including autonomous exploration, obstacle avoidance, and image-goal seeking, without requiring explicit global mapping. Code and model checkpoints are available at https://github.com/Zichen-Yan/SIGN.
@@ I Introduction
Learning vision-based navigation has attracted increasing attention in autonomous drone research. Existing works mostly focus on achieving autonomous obstacle avoidance and reference tracking, where the reference signal can be a target position [flying], velocity [hu2025seeing, bhattacharya2025vision, davide1] or direction [Navrl, davide3]. ImageNav poses a further challenge: enabling autonomous drones to search an unknown environment for a given goal image while avoiding collisions. This problem remains largely unexplored but has significant value in practical applications such as disaster rescue and industrial inspection, especially when pre-built maps and external localization are unavailable. Purely vision-based navigation reduces the dependence on additional sensors to enable a low-cost and lightweight drone design. In this paper, we investigate ImageNav for autonomous drones using only visual and IMU measurements.
The body of ImageNav research has primarily focused on ground robots [zer, fgp, ovrlv2, zson]. Most existing studies are conducted exclusively in simulators such as Habitat [habitat2] and neglect to penalize collisions with obstacles, which can result in unsafe behaviors when deployed in the real world. Meanwhile, discrete action spaces such as {move_forward, turn_left, turn_right, stop} are commonly adopted to simplify learning by reducing the search space. This is generally only acceptable for ground robots, which are inherently stable and can correct their position through vision by the discrete low-frequency actions.
In contrast, achieving safe and efficient ImageNav for autonomous drones is considerably more challenging. Drones are inherently unstable platforms with fast dynamics, necessitating high-frequency feedback control and accurate global localization to maintain stable flight. When relying on visual-inertial localization, they are particularly susceptible to pose-estimation drift, especially in indoor environments. Furthermore, discrete-action policies often lack the precision and responsiveness needed for rapid pose correction. For flight control systems, frequent switching between discrete actions can severely degrade real-time tracking performance due to the abrupt change of references. In this case, a continuous policy that directly outputs velocity commands can address this issue by enabling smoother and more timely control. Continuous actions also support seamless integration with flight controllers without additional planning modules. However, it also expands the exploration space, substantially increasing training complexity. In addition, naively embedding safety constraints into policy learning can hinder effective exploration and degrade navigation performance, particularly in cluttered or narrow environments such as Gibson [gibson], MP3D [mp3d], and HM3D [hm3d]. Therefore, balancing efficient exploration with robust real-world safety remains a central challenge in autonomous drone navigation.
In this paper, we introduce SIGN (Safety-aware Image-Goal Navigation), an end-to-end RL-based ImageNav framework for autonomous drones that employs continuous velocity control and incorporates a safety module for collision avoidance. Our main contributions are as follows:
@@ II-A Visual Navigation on Drones
We focus on end-to-end, map-less vision-based navigation without relying on global localization. Prior research in this direction has mainly addressed reference-tracking control with obstacle avoidance [flying, hu2025seeing, davide1, bhattacharya2025vision, Navrl] or drone racing on a fixed course [davide2], where an RL policy maps sensor inputs directly to control commands such as velocity or acceleration. The key distinction from these tasks is that ImageNav specifies the reference signal through an image, which serves as a visual goal for navigation. Similarly, Tao [tao2024learning] explored RL-based indoor search using drones, but their approach depends on global mapping. In contrast, our task is more challenging as it requires not only map-less exploration and collision avoidance, but also navigation toward a specified goal image.
@@ II-B End-to-end RL-based Visual Navigation
For visual navigation tasks, end-to-end RL has demonstrated higher efficiency in real-time inference and module tuning by avoiding intermediate stages such as mapping and subgoal selection [neuralslam]. Early work by Wijmans [ddppo] addressed point-goal visual navigation for ground robots. However, it relied on GPS and compass data, which is impractical for indoor settings. Al-Halah [zer] removed the dependency on localization and proposed a map-less visual navigation framework, enabling zero-shot generalization to new tasks and multi-modal goals. Additionally, previous works have employed memory mechanisms [memaug] and topological maps [vgm, tsgm, DBLP:conf/nips/HahnCTMRG21] to enhance exploration efficiency by storing keyframes and leveraging graph neural networks to extract visual correlations for guidance. For now, most existing end-to-end RL methods are constrained to discrete actions and suffer from training inefficiency due to environmental diversity. To better facilitate drone deployment and simplify the sim-to-real transfer, we investigate the extension from discrete to continuous settings.
@@ II-C Image-conditioned Waypoint Planner
Many works predict waypoints to reach a target image and use low-level position controllers for trajectory tracking [nomad, vint, gnm], even assisted by a world model [bar2025navigation]. A common weakness of these methods lies in the out-of-distribution (OOD) issues inherent to the supervised learning paradigm, especially when the current and target images have no overlap. In many image-conditioned navigation methods, continuous policies have already been explored through imitation learning. However, ImageNav involves numerous OOD scenarios that make generalization challenging without global localization to guide exploration. To address this issue, an RL-based continuous policy presents a promising solution. Additionally, an emerging trend is to combine 3D reconstruction techniques with ImageNav for drones, such as NeRF [adamkiewicz2022vision] and Gaussian Splatting [SINGER]. They typically require offline reconstruction and substantial computational resources. Due to the limited onboard capacity of drones, real-time processing of such methods remains impractical.
Figure 2: Performance comparison between discrete and continuous action using the SoTA baseline FGPrompt [fgp].
@@ III Visual Reinforcement Learning for ImageNav
As depicted in Fig. 1, SIGN is an end-to-end framework for drone ImageNav that leverages a continuous policy to enable map-less navigation with enhanced safety. But training continuous policies for ImageNav presents a significant challenge due to the expanded action space and increased training complexity. In Fig. 2, we compare a discrete State-of-The-Art (SoTA) baseline with its continuous counterpart, where the discrete action space $\mathcal{A}_{d}=\{\textit{move\_forward},\textit{turn\_left},\textit{turn\_right},\textit{stop}\}$ and the continuous one consists of forward linear and angular velocities, $\mathcal{A}_{c}=\{v_{lin},v_{ang}\}$. It can be observed that the discrete-action policy achieves steadily increasing rewards and much higher training efficiency, while the continuous-action policy suffers from very slow reward growth. This highlights the difficulty of training continuous policies in ImageNav tasks and motivates the design of our SIGN framework to address these challenges.
@@ III-B Navigation Model
The navigation model consists of a ResNet encoder $f_{\omega}$, a GRU model $g_{\xi}$, and an actor-critic policy trained by Proximal Policy Optimization (PPO) [ppo]. See Fig. 3 for the diagram. The goal image $O_{g}$ and current image $O_{c}$ are stacked as the input $s_{t}$. The inference procedure can be summarized as
@@ III-C Augmented Continuous Policy Training
A possible reason of the inefficient continuous policy training in Fig. 2 is that, in the early stage, training an effective vision encoder from scratch is time-consuming. RL policy struggles to improve until the vision encoder learns to capture key information from pixels. To address this problem, we propose two useful Self-Supervised Learning (SSL) techniques as auxiliary tasks to enhance the policy learning.
Recall the standard optimization form in vanilla PPO [ppo]
@@ III-C1 Future Prediction
The prediction ability receives widespread attention due to its success in the model-based visual RL research [dreamerv3, bar2025navigation], where a foundation world model acts as an internal simulator to imagine future observations and guide policy learning. Focusing on a portable model for drone applications, here we train a single-step prediction to enhance the visual backbone training. To this end, future prediction task is introduced, as illustrated in Fig. 4. We deploy an MLP head $P_{\psi}$ after the visual encoder $f_{\omega}$ to predict the reward $r_{t}$, the next-step latent feature $h_{t+1}=f_{\omega}(s_{t+1})$, and the done flag $d_{t}$. The auxiliary training loss is defined as
@@
The forward prediction of dynamics $\hat{h}_{t+1}=P_{\psi}(a_{t},h_{t})$ is implemented in the latent space to avoid deconvolution operations and save computation resources. The reward loss function uses cross-entropy between the predicted reward $\hat{r}_{t}$ and a two-hot encoding of reward $r_{t}$, which proves more effective in predicting sparse rewards and robust to reward magnitude [mrq].
@@ III-C2 RandomShift
The RandomShift [DBLP:conf/iclr/YaratsFLP22] technique performs data augmentation on the current image $O_{c}$ in each sample batch. Denote the perturbated data as $aug(s_{t})=randomshift(O_{c})\oplus O_{g}$, then auxiliary loss is defined as
@@
1:Input: number of env $B$, trajectory length $L$, fixed weights $\lambda_{r},\lambda_{d},\lambda_{T},\lambda_{rs}$
2:Initialize actor-critic models $A_{\theta}$ and $V_{\phi}$ including the backbone $f_{\omega}$ and GRU $g_{\xi}$
3:Initialize predictor model $P_{\psi}$ and collision predictor $Q_{c}$
4:for each epoch do
5:  Collect one trajectory $\{s_{t},a_{t},s_{t+1},r_{t},c_{t},d_{t}\}_{1:L}$ from each env, a total of $B\times L$ samples # Future Prediction
6:  for each mini-batch do
7:   Update $P_{\psi}$ and $f_{\omega}$ with data $\{s_{t},a_{t},s_{t+1},r_{t},d_{t}\}$ via minimizing Eq. (3)
8:  end for# PPO training with RandomShift
9:  for each mini-batch do
10:   Update policy parameters $\omega,\xi,\theta,\phi$ via Eq. (5)
11:   Train collision predictor $Q_{c}$ via minimizing Eq. (9)
12:  end for
13:end for
@@ IV Safety Module for Real-world Experiments
Existing ImageNav algorithms primarily focus on navigation performance while neglecting safety. Navigation policies trained in Habitat in those works are allowed to collide with obstacles. In practice, enforcing safety during policy training can lead to conservative behavior, which significantly hinders exploration in narrow indoor environments. Inspired by the safety filter designs in the fields of safe RL and control [usl, wabersich2021predictive, brunke2022safe], we propose a practical action correction mechanism that is activated only when the agent is in danger. This decoupled design is shown to be effective in balancing exploration and safety. Meanwhile, it also simplifies training.
@@ IV-A Model Training for Collision Probability Prediction
The depth image is first preprocessed by cropping and min-pooling operations to obtain a depth vector $s^{d}_{t}\in\mathbb{R}^{n}$, indicating the closest distance in different directions. $Q_{c}$ is used to predict the collision probability based on the current depth image and navigation action $a_{t}$, defined as
@@
where $c_{t}$ is a collision feedback from the Habitat simulator.
@@
$c_{t}^{\text{soft}}=\text{Sigmoid}\left(10\cdot(\beta-\min_{i}s^{d}_{t}(i))\right),$ (7)
where $\min(\cdot)$ denotes the minimum element of vector $s_{t}^{d}$ and $\beta$ is the threshold for a safe distance. When the minimum perceived distance falls below $\beta$, the soft label value increases rapidly toward 1, and it decreases to 0 when the distance exceeds $\beta$. Thus, collision detection is defined with respect to a safety margin distance of $\beta$.
@@ IV-B Action Correction for Obstacle Avoidance
The predicted collision probability is conditioned on $s_{t}^{d}$ and $a_{t}$, which means we can modify the action to reduce the collision probability. The correction mechanism is illustrated in Fig. 5. When the value of $Q_{c}$ exceeds a pre-defined threshold $d_{c}$, it indicates a dangerous situation, then actions from the navigation policy will be corrected iteratively until $Q_{c}<d_{c}$. The correction function $\varphi(\cdot)$ will project the raw action $a_{0}=a_{t}$ to a safer one based on gradients [usl]
$a_{k+1}=\varphi(a_{k})=a_{k}-\frac{\partial}{\partial a_{k}}[Q_{c}(s_{t}^{d},a_{k})-d_{c}]^{+},$ (10)
where $k=0,1,\ldots,n$ is the iteration number. The adjustment of the probability threshold $d_{c}$ can be viewed as a trade-off between safety and navigation performance. In general cases, $d_{c}$ is set to 0.5 because when $Q_{c}>0.5$, the minimum distance is less than the safety distance $\beta$ according to Eq. (7).
@@ V-A1 Datasets
Our navigation model is trained on Gibson ImageNav dataset [gibson], including 72 training scenes and 14 testing scenes. To further validate the robustness of our algorithm, the trained model is also evaluated on cross-domain datasets MP3D [mp3d] and HM3D [hm3d]. Scenes from different datasets have varying backgrounds and layouts but are all simulated in the Habitat.
@@ V-A2 Agent Configurations
To be consistent with SoTA works, we employ the Habitat camera that has a $90^{\circ}$ Field of View (FoV) and $128\times 128$ resolution. The continuous action space $\mathcal{A}_{c}$ consists of forward linear velocity $v_{lin}\in[0,0.25]~m/s$ and angular velocity $v_{ang}\in[-15,15]~^{\circ}/s$. To implement the ‘Stop’ action in a continuous setting, the episode will be terminated early when $v_{lin}<0.025~m/s\ \&\ |v_{ang}|<1.5~^{\circ}/s$.
@@
Mem-Aug [memaug] ResNet18 4 RGB Discrete 56.0% 69.0%
VGM [vgm] ResNet18 4 RGB Discrete 64.0% 76.0%
TSGM [tsgm] ResNet18 4 RGB Discrete 67.2% 81.1%
ZER [zer] ResNet9 RGB Discrete 21.6% 29.2%
ZSON [zson] ResNet50 RGB Discrete 28.0% 36.9%
OVRL-V2 [ovrlv2] ViT-Base RGB Discrete 58.7% 82.0%
FGPrompt [fgp] ResNet9 RGB Discrete 66.5% 90.4%
FGPrompt [fgp] ResNet9 RGB Continuous 35.0% 79.9%
@@
Mem-Aug [memaug] Discrete $3.9\%$ $6.9\%$ $3.5\%$ $1.9\%$
NRNS [DBLP:conf/nips/HahnCTMRG21] Discrete $5.2\%$ $9.3\%$ $4.3\%$ $6.6\%$
ZER [zer] Discrete $10.8\%$ $14.6\%$ $6.3\%$ $9.6\%$
FGPrompt [fgp] Discrete $50.4\%$ $77.6\%$ $49.6\%$ $76.1\%$
FGPrompt [fgp] Continuous $22.9\%$ $59.4\%$ $20.1\%$ $56.5\%$
@@
V-A3 Reward and Performance Metrics
The design of the reward function $r_{t}$ follows ZER [zer], which encourages the policy to approach the target and reduce the angle between the current view and the view of the goal.
In our experiments, $d_{s}=1~m$ and $\alpha_{s}=25^{\circ}$ denote the success thresholds for distance and angular deviation, respectively. $r_{suc}=2.5$ is the success reward and $\gamma=-0.01$ is the slack penalty. On benchmarks, Success Rate (SR) and Success Weighted by Path Length (SPL) are two main metrics to quantify the navigation performance. SPL is calculated as $\frac{1}{N}\Sigma^{N}_{i=1}S_{i}\frac{L_{i}}{max(L_{i},p_{i})}$, where $N$ is the total number of episodes, $S_{i}\in\{0,1\}$ is a success indicator, $p_{i}$ is the actual path length in the $i$-th episode, and $L_{i}$ is the optimal path length from the start to the goal point.